OpenStudy (anonymous):

Vong grilled 21 burgers at a block party. He grilled the same number of pounds of turkey burgers as hamburgers. Each turkey burger weighed 1/4 pound and each hamburger weighed 1/3 pound. How many of each did Vong grill?

OpenStudy (anonymous):

I have two equations t+h=21 and 1/4t = 1/3h

OpenStudy (anonymous):

I am just not sure if they are right

OpenStudy (anonymous):

If they are then t=12 and h=9

OpenStudy (anonymous):

okay, I checked, they work.

OpenStudy (anonymous):

Let t be the number of turkey burgers. From the problem statement: t (1/4) = (21 - t) (1/3) t = 12